# On the Embedding Collapse When Scaling up Recommendation Models

### v1

@@ Appendix B Non-Linearity for Multi-Embedding
By simply letting ${\bm{W}}^{(m,m)}={\bm{W}}{(m)}$ and ${\bm{W}}^{(m,m^{\prime})}={\bm{O}}$ for $m\neq m^{\prime}$, we convert a multi-embedding model into a single-embedding model under permutation. Therefore, multi-embedding is a special case of single-embedding for DCNv2.
@@
Appendix C Details of Experiment
C.1 Dataset Description
Table 2: Statistics of benchmark datasets for experiments.
Dataset #Instances #Fields $\sum_{i}D_{i}$ $\max\{D_{i}\}$ $\min\{D_{i}\}$
Criteo 45.8M 39 1.08M 0.19M 4
Avazu 40.4M 22 2.02M 1.61M 5
C.2 Experiment Settings
Specific multi-embedding design.
Hyperparameters.
C.3 Experimental Results
Table 3: Results on Criteo dataset. Higher indicates better.
Model Criteo
base 2x 3x 4x 10x
DNN SE 0.81228±${}_{\pm 0.00004}$ 0.81222±${}_{\pm 0.00002}$ 0.81207±${}_{\pm 0.00007}$ 0.81213±${}_{\pm 0.00011}$ 0.81142±${}_{\pm 0.00006}$
ME 0.81261±${}_{\pm 0.00004}$ 0.81288±${}_{\pm 0.00015}$ 0.81289±${}_{\pm 0.00007}$ 0.81287±${}_{\pm 0.00005}$
IPNN SE 0.81272±${}_{\pm 0.00003}$ 0.81273±${}_{\pm 0.00013}$ 0.81272±${}_{\pm 0.00004}$ 0.81271±${}_{\pm 0.00007}$ 0.81262±${}_{\pm 0.00016}$
ME 0.81268±${}_{\pm 0.00009}$ 0.81270±${}_{\pm 0.00002}$ 0.81273±${}_{\pm 0.00015}$ 0.81311±${}_{\pm 0.00008}$
NFwFM SE 0.81059±${}_{\pm 0.00012}$ 0.81087±${}_{\pm 0.00008}$ 0.81090±${}_{\pm 0.00012}$ 0.81112±${}_{\pm 0.00011}$ 0.81113±${}_{\pm 0.00022}$
ME 0.81128±${}_{\pm 0.00017}$ 0.81153±${}_{\pm 0.00002}$ 0.81171±${}_{\pm 0.00012}$ 0.81210±${}_{\pm 0.00010}$
xDeepFM SE 0.81217±${}_{\pm 0.00003}$ 0.81180±${}_{\pm 0.00002}$ 0.81167±${}_{\pm 0.00008}$ 0.81137±${}_{\pm 0.00005}$ 0.81116±${}_{\pm 0.00009}$
ME 0.81236±${}_{\pm 0.00006}$ 0.81239±${}_{\pm 0.00022}$ 0.81255±${}_{\pm 0.00011}$ 0.81299±${}_{\pm 0.00009}$
DCNv2 SE 0.81339±${}_{\pm 0.00002}$ 0.81341±${}_{\pm 0.00007}$ 0.81345±${}_{\pm 0.00009}$ 0.81346±${}_{\pm 0.00011}$ 0.81357±${}_{\pm 0.00004}$
ME 0.81348±${}_{\pm 0.00005}$ 0.81361±${}_{\pm 0.00014}$ 0.81382±${}_{\pm 0.00015}$ 0.81385±${}_{\pm 0.00005}$
FinalMLP SE 0.81259±${}_{\pm 0.00009}$ 0.81262±${}_{\pm 0.00007}$ 0.81248±${}_{\pm 0.00008}$ 0.81240±${}_{\pm 0.00002}$ 0.81175±${}_{\pm 0.00020}$
ME 0.81290±${}_{\pm 0.00017}$ 0.81302±${}_{\pm 0.00005}$ 0.81303±${}_{\pm 0.00004}$ 0.81303±${}_{\pm 0.00012}$
Figure 10: Visualization of scalability on Criteo dataset.
Table 4: Results on Avazu dataset. Higher indicates better.
Model Avazu
base 2x 3x 4x 10x
DNN SE 0.78744±${}_{\pm 0.00008}$ 0.78759±${}_{\pm 0.00011}$ 0.78752±${}_{\pm 0.00031}$ 0.78728±${}_{\pm 0.00036}$ 0.78648±${}_{\pm 0.00013}$
ME 0.78805±${}_{\pm 0.00017}$ 0.78826±${}_{\pm 0.00013}$ 0.78862±${}_{\pm 0.00026}$ 0.78884±${}_{\pm 0.00005}$
IPNN SE 0.78732±${}_{\pm 0.00020}$ 0.78741±${}_{\pm 0.00022}$ 0.78738±${}_{\pm 0.00010}$ 0.78750±${}_{\pm 0.00007}$ 0.78745±${}_{\pm 0.00018}$
ME 0.78806±${}_{\pm 0.00012}$ 0.78868±${}_{\pm 0.00023}$ 0.78902±${}_{\pm 0.00009}$ 0.78949±${}_{\pm 0.00028}$
NFwFM SE 0.78684±${}_{\pm 0.00017}$ 0.78757±${}_{\pm 0.00020}$ 0.78783±${}_{\pm 0.00009}$ 0.78794±${}_{\pm 0.00022}$
ME 0.78868±${}_{\pm 0.00038}$ 0.78901±${}_{\pm 0.00029}$ 0.78932±${}_{\pm 0.00035}$
xDeepFM SE 0.78743±${}_{\pm 0.00009}$ 0.78750±${}_{\pm 0.00025}$ 0.78714±${}_{\pm 0.00030}$ 0.78735±${}_{\pm 0.00004}$ 0.78693±${}_{\pm 0.00050}$
ME 0.78848±${}_{\pm 0.00006}$ 0.78886±${}_{\pm 0.00026}$ 0.78894±${}_{\pm 0.00004}$ 0.78927±${}_{\pm 0.00019}$
DCNv2 SE 0.78786±${}_{\pm 0.00022}$ 0.78835±${}_{\pm 0.00023}$ 0.78854±${}_{\pm 0.00010}$ 0.78852±${}_{\pm 0.00003}$ 0.78856±${}_{\pm 0.00016}$
ME 0.78862±${}_{\pm 0.00011}$ 0.78882±${}_{\pm 0.00012}$ 0.78907±${}_{\pm 0.00011}$ 0.78942±${}_{\pm 0.00024}$
FinalMLP SE 0.78751±${}_{\pm 0.00026}$ 0.78797±${}_{\pm 0.00019}$ 0.78795±${}_{\pm 0.00017}$ 0.78742±${}_{\pm 0.00015}$ 0.78662±${}_{\pm 0.00025}$
ME 0.78821±${}_{\pm 0.00013}$ 0.78831±${}_{\pm 0.00029}$ 0.78836±${}_{\pm 0.00018}$ 0.78830±${}_{\pm 0.00022}$
Figure 11: Visualization of scalability on Avazu dataset.

### v2

@@ Appendix E Non-Linearity for Multi-Embedding
By simply letting ${\bm{W}}^{(m,m)}={\bm{W}}^{(m)}$ and ${\bm{W}}^{(m,m^{\prime})}={\bm{O}}$ for $m\neq m^{\prime}$, we convert a multi-embedding model into a single-embedding model under permutation. Therefore, multi-embedding is a special case of single-embedding for DCNv2.
@@
Appendix F Detailed Explanation of Embedding Diversity
In Section 5.3, we propose to use principal angle to measure embedding set diversity. Here we introduce the motivation and an example. Note that
$\mathrm{rank}\left(\left[{\bm{E}}^{(m)},{\bm{E}}^{(m^{\prime})}\right]\right)$ $=\mathrm{rank}\left(\left[{\bm{U}}^{(m)}{\bm{\Sigma}}^{(m)}\big{(}{\bm{V}}^{(m)}\big{)}^{\top},{\bm{U}}^{(m^{\prime})}{\bm{\Sigma}}^{(m^{\prime})}\big{(}{\bm{V}}^{(m^{\prime})}\big{)}^{\top}\right]\right)$
$=\mathrm{rank}\left(\left[{\bm{U}}^{(m)}{\bm{\Sigma}}^{(m)}\big{(}{\bm{V}}^{(m)}\big{)}^{\top},{\bm{U}}^{(m^{\prime})}{\bm{\Sigma}}^{(m^{\prime})}\big{(}{\bm{V}}^{(m^{\prime})}\big{)}^{\top}\right]\begin{bmatrix}{\bm{V}}^{(m)}&{\bm{O}}\\ {\bm{O}}&{\bm{V}}^{(m^{\prime})}\end{bmatrix}\right)$
$=\mathrm{rank}\left(\left[{\bm{U}}^{(m)}{\bm{\Sigma}}^{(m)},{\bm{U}}^{(m^{\prime})}{\bm{\Sigma}}^{(m^{\prime})}\right]\right)$
$=\mathrm{rank}\left({\bm{U}}^{(m)}{\bm{\Sigma}}^{(m)}\right)+\mathrm{rank}\left({\bm{U}}^{(m^{\prime})}{\bm{\Sigma}}^{(m^{\prime})}\right)-\mathrm{rank}\left(\left({\bm{U}}^{(m)}{\bm{\Sigma}}^{(m)}\right)^{\top}{\bm{U}}^{(m^{\prime})}{\bm{\Sigma}}^{(m^{\prime})}\right)$
$=\mathrm{rank}\left({\bm{E}}^{(m)}\right)+\mathrm{rank}\left({\bm{E}}^{(m^{\prime})}\right)-\mathrm{rank}\left(\big{(}{\bm{U}}^{(m)}\big{)}^{\top}{\bm{U}}^{(m^{\prime})}\right),$
where the second last line are derived from the orthonormality of ${\bm{U}}$. Note that
$\mathrm{rank}\left(\big{(}{\bm{U}}^{(m)}\big{)}^{\top}{\bm{U}}^{(m^{\prime})}\right)=\|\cos\left(\bm{\phi}^{m\leftrightarrow m^{\prime}}\right)\|_{0},$
and we therefore generalize it to $\frac{1}{K}\|\cos\left(\bm{\phi}^{m\leftrightarrow m^{\prime}}\right)\|_{1}$ to measure similarity, and use $1-\mathrm{similarity}$ as diversity.
Considering the following example of diversity. An embedding with size of $2$ are learned as
${\bm{E}}=\begin{pmatrix}1&0\\ 1&0\\ 0&1\\ 0&1\end{pmatrix}$
with $\mathrm{rank}({\bm{E}})=\mathrm{IA}({\bm{E}})=2$. If enlarged to size of $4$, due to interaction-collapse theory, it is likely to be learned as
${\bm{E}}^{(1)}=\begin{pmatrix}1&0\\ 1&0\\ 0&1\\ 0&1\end{pmatrix},\quad{\bm{E}}^{(2)}=\begin{pmatrix}1&0\\ 1&0\\ 0&1\\ 0&1\end{pmatrix},\quad{\bm{E}}^{\text{single}}=\begin{pmatrix}1&0&1&0\\ 1&0&1&0\\ 0&1&0&1\\ 0&1&0&1\end{pmatrix},$
with $\cos\left(\bm{\phi}^{1\leftrightarrow 2}\right)=(1,1)$, $\mathrm{rank}({\bm{E}}^{\text{single}})=\mathrm{IA}({\bm{E}}^{\text{single}})=2$, i.e., the enlarged size does not increase information abundance. Here
When using multi-embedding, the embedding sets are possibly learned to be of large diversity, and the overall embedding are learned as
${\bm{E}}^{(1)}=\begin{pmatrix}1&0\\ 1&0\\ 0&1\\ 0&1\end{pmatrix},\quad{\bm{E}}^{(2)}=\begin{pmatrix}1&0\\ 0&1\\ 1&0\\ 0&1\end{pmatrix},\quad{\bm{E}}^{\text{multi}}=\begin{pmatrix}1&0&1&0\\ 1&0&0&1\\ 0&1&1&0\\ 0&1&0&1\end{pmatrix},$
with $\cos\left(\bm{\phi}^{1\leftrightarrow 2}\right)=(1,0)$, $\mathrm{rank}({\bm{E}}^{\text{multi}})=3$ and $\mathrm{IA}({\bm{E}}^{\text{multi}})=1+\sqrt{2}$, indicating the effectiveness of multi-embedding.
Appendix G Details of Toy Expermient
Appendix H Empirical analysis on FFM in Evidence I
where sub-embedding ${\bm{E}}_{i}^{\to j}\in\mathbb{R}^{D_{i}\times K/(N-1)}$ is only used when interacting field $i$ with field $j$ for $j\neq i$. We perform the same experiments as Evidence I, and similarly find that $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ are co-influenced by both $\mathrm{IA}({\bm{E}}_{i})$ and $\mathrm{IA}({\bm{E}}_{j})$, as shown in Figure 14. This is amazing in the sense that even using independent embeddings to represent the same field features, these embeddings get different information abundance after learning.
(a) $\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
(b) $\sum\limits_{j=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
(c) $\sum\limits_{i=1}^{N}\mathrm{IA}({\bm{E}}_{i}^{\to j})$.
Figure 14: Information abundance of sub-embedding matrices for FFM, with field indices sorted by information abundance of corresponding raw embedding matrices. Higher or warmer indicates larger. Similarly, $\mathrm{IA}({\bm{E}}_{i}^{\to j})$ are co-influenced by both $\mathrm{IA}({\bm{E}}_{i})$ and $\mathrm{IA}({\bm{E}}_{j})$.
Appendix I Extension of Information Abundance
[…]
Figure 15: Normalized information abundance $\frac{\mathrm{IA}({\bm{E}})}{K}$ for different embedding sizes on DCNv2.
Appendix J Detailed Explanation of Regularized DCNv2
Regarding Evidence II, we proposed regularization of the weight matrix ${\bm{W}}_{i}^{\to j}$ to mitigate the collapse caused by the projection ${\bm{W}}_{i}^{\to j}$ in sub-embeddings. By regularizing ${\bm{W}}_{i}^{\to j}$ to be a unitary matrix (or the multiplication of unitary matrices), we ensure the preservation of all singular values of the sub-embedding. Consequently, the information abundance of sub-embeddings in regularized DCNv2 is larger than standard DCNv2. We plot the heatmap of information abundance of embeddings and sub-embeddings Figure 16. This clearly demonstrates that regularized DCNv2 exhibits a higher information abundance. Based on our Finding 1, regularized DCNv2 mitigates the problem of embedding collapse by increasing the information abundance of the sub-embeddings that directly interact with the embeddings.
(a) Standard DCNv2
(b) Regularized DCNv2
Figure 16: Visualization of information abundance of embeddings and sub-embeddings for standard and regularized DCNv2, respectively. The rightmost and downmost bars correspond to $\mathrm{IA}({\bm{E}}_{i})$ or $\mathrm{IA}({\bm{E}}_{j})$. Compared with the standard DCNv2, the regularized can preserve singular values, and resulting in less-collapsed sub-embeddings and finally larger information abundance as in Figure 5a.
Appendix K How ME Performs When Feature Interaction Is Suppressed?
In this section we analysis of ME under models with feature interaction suppressed, as discussed Section 4.2, where SE suffers from overfiting.
Evidence III for ME.
We add regularization
$\ell_{reg}=\sum_{m=1}^{M}\sum_{i=1}^{N}\sum_{j=1}^{N}\left\|({\bm{W}}^{(m)}_{i\to j})^{\top}{\bm{W}}^{(m)}_{i\to j}-\lambda^{(m)}_{ij}{\bm{I}}\right\|_{\mathrm{F}}^{2}$
for ME DCNv2 and conduct experiments with different embedding sizes. Results are shown in Figure 17a. Even though the performance is worse than that without regularization, compared with SE, ME still consistently improves the performance with the growth of model size.
(a) Standard vs Regularized
(b) DCNv2 vs DNN
Figure 17: Test AUC w.r.t. model size.
Evidence IV for ME.
We compare the performance of SE/ME on DNN/DCNv2, as shown in Figure 17b. Compared with SE, ME DNN improves the performance with the growth of model size.
Summary: ME offers scalability even though feature interaction is suppressed.
For models with feature interaction suppressed such as regularized DCNv2 and DNN, the performance of SE might dropped when enlarging models, since feature interaction provides domain knowledge and large models might suffer from overfitting. Experiments show that these models with ME can properly scale up. Such results are plausible since ME improves scalability by capturing diverse patterns instead of learning with a single interaction pattern.